# New Taipei City, 251301, Taiwan

### v1

@@ 3.2 Feature Contexts Excitation Modules
Inference
Time
YOLOv8+GE M1 Small 640 11.28M 29.1G 62.44% 2.6ms
Medium 640 26.10M 79.8G 63.00% 5.1ms
Large 640 43.91M 166.3G 64.15% 7.6ms
YOLOv8+GE M2 Small 640 11.15M 28.7G 62.07% 2.6ms
Medium 640 25.87M 79.1G 63.20% 5.1ms
Large 640 43.65M 165.4G 64.02% 7.6ms
YOLOv8+GE M3 Small 640 11.16M 28.7G 61.89% 2.7ms
Medium 640 25.90M 79.1G 62.22% 5.2ms
Large 640 43.68M 165.5G 62.42% 7.7ms
YOLOv8+GE M1 Small 1024 11.28M 29.1G 63.72% 5.2ms
Medium 1024 26.10M 79.8G 63.81% 10.0ms
Large 1024 43.91M 166.3G 64.34% 15.4ms
YOLOv8+GE M2 Small 1024 11.15M 28.7G 62.51% 5.1ms
Medium 1024 25.87M 79.1G 64.28% 10.0ms
Large 1024 43.65M 165.4G 65.99% 15.3ms
YOLOv8+GE M3 Small 1024 11.16M 28.7G 62.42% 4.8ms
Medium 1024 25.90M 79.1G 62.63% 10.3ms
Large 1024 43.68M 165.5G 64.35% 15.7ms
Table 4: Experimental results of selecting improved methods for pediatric wrist fracture detection using the GRAZPEDWRI-DX dataset for YOLOv8+GCT model. The best performance and the 2nd best performance are in red and blue colors, respectively.
Model Method
Model
Size
Input
Size
@@
Inference
Time
YOLOv8+GCT M1 Small 640 11.27M 29.1G 61.91% 2.7ms
Medium 640 26.08M 79.8G 62.64% 5.1ms
Large 640 43.90M 166.3G 63.70% 7.6ms
YOLOv8+GCT M2 Small 640 11.14M 28.7G 61.36% 2.6ms
Medium 640 25.86M 79.1G 62.19% 5.1ms
Large 640 43.64M 165.4G 63.42% 7.6ms
YOLOv8+GCT M3 Small 640 11.14M 28.7G 52.07% 2.7ms
Medium 640 25.86M 79.1G 53.60% 5.2ms
Large 640 43.64M 165.4G 57.88% 7.6ms
YOLOv8+GCT M1 Small 1024 11.27M 29.1G 62.49% 5.1ms
Medium 1024 26.08M 79.8G 63.47% 10.1ms
Large 1024 43.90M 166.3G 63.87% 15.3ms
YOLOv8+GCT M2 Small 1024 11.14M 28.7G 63.51% 5.0ms
Medium 1024 25.86M 79.1G 64.46% 10.1ms
Large 1024 43.64M 165.4G 65.67% 15.3ms
YOLOv8+GCT M3 Small 1024 11.14M 28.7G 52.53% 5.1ms
Medium 1024 25.86M 79.1G 53.70% 10.1ms
Large 1024 43.64M 165.4G 58.15% 15.7ms
Table 5: Quantitative comparison (Params/FLOPs/F1-Score/mAP@50/mAP@50-95/Inference time per image) with other SOTA models for pediatric wrist fracture detection on the GRAZPEDWRI-DX dataset. The sizes of YOLOv8 and its improved models are all large (L), and the size of YOLOv9 model is extended (E). The best performance and the 2nd best performance are in red and blue colors, respectively.
Model
Input
Size
@@
Figure 6: Visualization of the accuracy of predicting each class using our proposed FCE-YOLOv8 models and other SOTA models on the GRAZPEDWRI-DX dataset with the input image size of 1024: (a) YOLOv8+ResCBAM, (b) YOLOv9, (c) YOLOv8+GC, (d) YOLOv8+GCT, (e) YOLOv8+SE, and (f) YOLOv8+GE.
@@
4 Experiments
4.1 Datasets
4.2 Evaluation Metrics
Fracture detection on pediatric wrist X-ray images belongs to object detection tasks, and six evaluation metrics in object detection tasks are applicable, including the model parameters (Params), floating-point operations (FLOPs), F1-Score, mean average precision at 50% intersection over union (mAP@50), mean average precision from 50% to 95% intersection over union (mAP@50-95), and the inference time per image. Therefore, this work compares different model performance using these six evaluation metrics.
4.2.1 Params
in the model are determined by the architecture layer count, the number of neurons per layer, and the overall complexity of the model. In general, an increase in Params corresponds to a larger model size. While larger models generally offer better model performance, they also require more computational resources. Therefore, it is important to balance the model size with computational cost in practical applications.
4.2.2 FLOPs
representing floating-point operations, are commonly used to evaluate the computational complexity of the model. This metric serves as an indicator of the computational performance and speed of the models. In general, the models with lower FLOPs are more suitable for resource-limited environments, while the models with higher FLOPs require more powerful hardware support.
4.2.3 F1-Score
is the harmonic mean of the Precision and Recall of the model, ranging from 0 to 1. The value closer to 1 indicates that the model has achieved a better balance between Precision and Recall. Conversely, if either Precision or Recall is close to 0, F1-Score will also approach 0, signaling poor model performance.
4.2.4 mAP
is a widely used metric for evaluating the performance of object detection models. In object detection tasks, the model aims to recognize objects within an image and determines their locations. For each class, the model calculates the area under the Precision-Recall curve, known as average precision (AP), which measures the model detection performance for that specific class. The mAP score is calculated by averaging the AP values across all classes, providing an overall measure of the model detection accuracy.
4.3 Data Preparation
To ensure fairness in model performance comparison, the same data is used for all model training. We randomly divide the GRAZPEDWRI-DX [33] dataset into training, validation, and test sets in the ratio of 70%, 20%, and 10%, respectively. Specifically, the training set contains 14,204 images (69.88%); the validation set contains 4,094 images (20.14%), and the test set contains 2,029 images (9.98%).
Furthermore, data augmentation is performed on the training set before all models training. Specifically, this work adjusts the contrast and brightness of the X-ray images using the addWeighted function in the Open Source Computer Vision Library (OpenCV), and expanding the original training dataset from 14,204 images to 28,408 images.
4.4 Training
4.5 Selection of Improved Methods
This work presents three different methods for improving YOLOv8, resulting in FCE-YOLOv8 models. Given four proposed FCE modules (i.e., SE, GC, GE, and GCT), we conduct the experiment to determine the most suitable improved method for each module. The experimental results are shown in Tables 123, and 4. As described in Section 3.1, we apply three methods (i.e., M1, M2, and M3) to train FCE-YOLOv8 with SE, GC, GE, and GCT modules, respectively. Apart from the “Method”, “Model Size”, and “Input Size” parameters, all other settings remain the same for all models.
Table 1 presents the experimental results for YOLOv8+SE. It can be seen that for the SE module, when the input image size is 640, choosing M2 to construct the YOLOv8+SE model can get the highest mAP@50 value of 64.08%; while the input image size is 1024, choosing M3 to construct the YOLOv8+SE model can get the highest mAP@50 value of 67.07%. In addition, with the input size of 640, Tables 23, and 4 show that the most suitable method for GC, GE, and GCT modules is all M1. YOLOv8+GC, YOLOv8+GE, and YOLOv8+GCT with M1 achieve the highest mAP@50 values of 64.55%, 64.15%, and 63.70%, respectively. However, when the input size is set to 1024, the most effective method for YOLOv8+GC is M3, resulting in the mAP@50 value of 66.32%. For both YOLOv8+GE and YOLOv8+GCT, M2 is the most effective, obtaining the mAP@50 values of 65.99% and 65.67%, respectively. Since the input image size is set to 1024 for comparisons with the SOTA models, as shown in Section 4.6, we select M3, M3, M2, and M2 for SE, GC, GE, and GCT, respectively, to construct FCE-YOLOv8.
@@ 7 Declarations
7.1 Author contributions statement
R.J. conceived and conducted the experiments, and wrote the main manuscript; C.C. contributed expertise on fractures and assisted in analyzing the results; E.X. performed the formal analysis and visualization; J.C. was responsible for funding acquisition, supervision, and manuscript editing. All authors have reviewed the manuscript.
7.2 Funding
@@
7.3 Competing interests
The authors have no financial or proprietary interests in any material discussed in this article.
7.4 Ethics approval
This research does not involve human participants and/or animals.

### v2

@@ 4.4 Training
GE-M1-S 11.28M 29.1G 62.44% 2.6ms
GE-M1-M 26.10M 79.8G 63.00% 5.1ms
GE-M1-L 43.91M 166.3G 64.15% 7.6ms
GE-M2-S 11.15M 28.7G 62.07% 2.6ms
GE-M2-M 25.87M 79.1G 63.20% 5.1ms
GE-M2-L 43.65M 165.4G 64.02% 7.6ms
GE-M3-S 11.16M 28.7G 61.89% 2.7ms
GE-M3-M 25.90M 79.1G 62.22% 5.2ms
GE-M3-L 43.68M 165.5G 62.42% 7.7ms
GE-M1-S 11.28M 29.1G 63.72% 5.2ms
GE-M1-M 26.10M 79.8G 63.81% 10.0ms
GE-M1-L 43.91M 166.3G 64.34% 15.4ms
GE-M2-S 11.15M 28.7G 62.51% 5.1ms
GE-M2-M 25.87M 79.1G 64.28% 10.0ms
GE-M2-L 43.65M 165.4G 65.99% 15.3ms
GE-M3-S 11.16M 28.7G 62.42% 4.8ms
GE-M3-M 25.90M 79.1G 62.63% 10.3ms
GE-M3-L 43.68M 165.5G 64.35% 15.7ms
4.5 Selection of Improved Methods
This work presents three different methods for improving model architecture, resulting in FCE-YOLOv8 models. Given four proposed FCE modules (i.e., SE, GC, GE, and GCT), we conduct the experiment to determine the most suitable improved method for each module. The experimental results are shown in Tables 123, and 4. As described in Section 3.1, we apply three methods (i.e., M1, M2, and M3) to train FCE-YOLOv8 with SE, GC, GE, and GCT modules, respectively. Apart from the parameters of module, model size, and input image size, all other settings remain the same for all models.
Table 1 presents the experimental results for YOLOv8+SE. It can be seen that for the SE module, when the input image size is 640, choosing M2 to construct the YOLOv8+SE model can get the highest mAP@50 value of 64.08%; while the input image size is 1024, choosing M3 to construct the YOLOv8+SE model can get the highest mAP@50 value of 67.07%. In addition, with the input size of 640, Tables 23, and 4 show that the most suitable method for GC, GE, and GCT modules is all M1. YOLOv8+GC, YOLOv8+GE, and YOLOv8+GCT with M1 achieve the highest mAP@50 values of 64.55%, 64.15%, and 63.70%, respectively. However, when the input size is set to 1024, the most effective method for YOLOv8+GC is M3, resulting in the mAP@50 value of 66.32%. For both YOLOv8+GE and YOLOv8+GCT, M2 is the most effective, obtaining the mAP@50 values of 65.99% and 65.67%, respectively. Since the input image size is set to 1024 for comparisons with the SOTA models, as shown in Section 4.6, we select M3, M3, M2, and M2 for SE, GC, GE, and GCT, respectively.
Table 4: Experimental results of GCT module based on YOLOv8 on the GRAZPEDWRI-DX dataset. 1The input image size is 640. 2The input image size is 1024.
@@ 4.5 Selection of Improved Methods
GCT-M1-S1 11.27M 29.1G 61.91% 2.7ms
GCT-M1-M1 26.08M 79.8G 62.64% 5.1ms
GCT-M1-L1 43.90M 166.3G 63.70% 7.6ms
GCT-M2-S1 11.14M 28.7G 61.36% 2.6ms
GCT-M2-M1 25.86M 79.1G 62.19% 5.1ms
GCT-M2-L1 43.64M 165.4G 63.42% 7.6ms
GCT-M3-S1 11.14M 28.7G 52.07% 2.7ms
GCT-M3-M1 25.86M 79.1G 53.60% 5.2ms
GCT-M3-L1 43.64M 165.4G 57.88% 7.6ms
GCT-M1-S2 11.27M 29.1G 62.49% 5.1ms
GCT-M1-M2 26.08M 79.8G 63.47% 10.1ms
GCT-M1-L2 43.90M 166.3G 63.87% 15.3ms
GCT-M2-S2 11.14M 28.7G 63.51% 5.0ms
GCT-M2-M2 25.86M 79.1G 64.46% 10.1ms
GCT-M2-L2 43.64M 165.4G 65.67% 15.3ms
GCT-M3-S2 11.14M 28.7G 52.53% 5.1ms
GCT-M3-M2 25.86M 79.1G 53.70% 10.1ms
GCT-M3-L2 43.64M 165.4G 58.15% 15.7ms
Table 5: Quantitative comparison with other models for pediatric wrist fracture detection on the GRAZPEDWRI-DX dataset. The sizes of YOLOv8 and its improved models are all large (L), and the size of YOLOv9 model is extended (E). The best performance and the 2nd best performance are in red and blue colors, respectively.
@@
Figure 6: Visualization of the accuracy of predicting each class using our proposed FCE-YOLOv8 models and other SOTA models on the GRAZPEDWRI-DX dataset with the input image size of 1024.
@@ 7 Declarations
This paper is an expanded paper from International Symposium on Intelligent Signal Processing and Communication Systems (ISPACS) held on December 10-13, 2024 in Kaohsiung, Taiwan.